OpenStudy (anonymous):

simplify the expression: -5+i/2i

OpenStudy (anonymous):

\[\frac{5+i}{2i}\times \frac{i}{i}\] is a good start to rewrite in standard form

OpenStudy (anonymous):

so then you get 5i^2/2i^2 = 5/2?

OpenStudy (anonymous):

-5/2

OpenStudy (anonymous):

\[\frac{-5+i}{2i}\times \frac{i}{i}=\frac{-5i-1}{-2}\] \[=\frac{1}{2}+\frac{5}{2}i\] in standard form

OpenStudy (anonymous):

oh okay thank you I multiplied it wrong

OpenStudy (anonymous):

yw